OpenStudy (anonymous):

Find dy/dx if y=sin^2 x

OpenStudy (raden):

use the chain rule

OpenStudy (raden):

y = sin^2 x y ' = 2 sin(x) (sin(x))' = 2 sin(x) cos(x) it could be simplied again, by using the identity : sin(2x) = 2 sin(x)cos(x) so, y ' = 2 sin(x) cos(x) = sin(2x)

OpenStudy (raden):

does that make sense ?

OpenStudy (zehanz):

Alternatively, you could write y=sin^2(x) as y=sin(x)*sin(x) and use the Product Rule: y' = cos(x)*sin(x) + sin(x)*cos(x) = 2sin(x)cos(x) = sin(2x), which is (of course) the same answer as RadEn's!